There's mom, dad, brother, and me. The sum of our ages is 83. 6 times dad equals 7 times mom and she is 3 times me.
Answer by stanbon(75887)

There's mom, dad, brother, and me. The sum of our ages is 83. 6 times dad equals 7 times mom and she is 3 times me.
------
Equations:
m + d + b + u = 83
6d = 7m
m = 3u
--------
Substitute for "d" and "u":
m + (7/6)m + b + (1/3)m = 83
-------
2m + 7m + 6b + 2m = 6*83
-------
11m + 6b = 6*83
======
Comment: You need one more equation relating m and b to solve this equation.
-------
Cheers,
Stan H.
=======